OpenStudy (anonymous):

Nina is bouncing on a trampoline her fourth bounce is 6ft high. Each bounce afterward is two-thirds high as the bounce before it. How high will her sixth bounce be.

OpenStudy (anonymous):

\[6\times \frac{2}{3}\times \frac{2}{3}= \frac{8}{3}\]